NASA ISS On-Orbit Status 28 January 2014

©NASA

NASA ISS On-Orbit Status 28 January 2014.

Today: Columbus Cycle 14 Transition: Col-CC operations to transition the ground and onboard Columbus systems from software version 13 to 14 is proceeding nominally.

Columbus cycle 14 made improvements to the system stability and robustness, incorporated provisions for Solid State Recorder (SSD) coming up on ATV-5, the Video Data Processing Unit (VDPU) MkII, new external payload High Definition Earth Viewing (HDEV) arriving on SpX-3 and RapidSCAT on SpX-4.

Story Time From Space: FE-3 Hopkins gathered materials, set up a camcorder and read the children's book Max Goes to the Space Station and Max Goes to the Moon. This is the second and third book readings for Story Time From Space, which combines science literacy outreach with simple demonstrations recorded aboard the ISS. Crewmembers will read a total of five science, technology, engineering and mathematics-related children's books in orbit and complete simple science concept experiments. Video and data collected during the demonstrations are downlinked to the ground and posted in a video library with accompanying educational materials on the Center for the Advancement of Science in Space (CASIS) Academy educational website.

ISS SERVIR Environmental Research and Visualization System (ISERV) Preparation: A new pointing mount for ISERV was launched on Orb-1 that will replace a failed pointing mount. Today, Hopkins and FE-6 Wakata reviewed the procedures to partially remove the ISERV hardware from WORF, detach the failed pointing mount from the telescope tube, attach the new pointing mount, and then reinstall. Following the procedure review Hopkins and Wakata had a conference with the Principal Investigator on the ground to discuss details of the procedure and to answer any questions. ISERV automatically takes images of Earth through a small telescope with an off-the-shelf digital camera mounted in the US Laboratory module. There is a significant and persistent need for image data to support environmental decision-making, and to assess and monitor the impact of disasters and other significant events of the Earth surface. The unique attributes of the ISS, both in terms of orbital characteristics and available human and infrastructure support, make it a highly desirable platform from which to acquire these data.

Extravehicular Activity (EVA) Tool and Hardware Transfer: FE-5 Mastracchio and Commander Kotov transfered and stowed USOS EVA tools and cameras that were used during yesterday's RS EVA 37A.

Waste & Hygiene Compartment (WHC) Water Container Remove and Replace (R&R): Mastracchio performed routine maintenance on the WHC by R&R of the Water Container [ЕДВ]. Following the R&R he manually filled the Flush Water Tank using the Hydraulic Interface Control Panel.

USOS Food Warmer Troubleshooting - Part 2: In an effort to gain confidence in the Express Rack 6 Locker 8 location, the crew moved the deployed spare food warmer S/N 1004 from the PS-28 strip plugged into Express Rack 6 Locker 3 to Express Rack 6 Locker 8. Food Warmer S/N 1001, which has been tripping the in the locker 8 location, was disconnected for now.

Today's Planned Tasks

ENERGY Diet Log of Breakfast [Complete]
Acoustic Dosimeter Setup [Complete]
Elektron-VM Liquid Unit Pressurization before Activation [Complete]
Removal of US TV cameras, Lights, and REBA power unit from Orlan-MK Suits [Complete]
Periodic Fitness Evaluation - Nom Ops [Complete]
WRS Water Sample Analysis [Complete]
HAM Radio Power Up [Complete]
US Tools and Equipment Handover from RS to USOS after EVA-37A [Complete]
TERMO-LADA. Visual Inspection of Hardware Operation [Complete]
Photo (EVA) - camera handover to USOS [Complete]
Refilling of Orlan No. 4, No. 6 water tank [Complete]
EVA-37A Photo Downlink and D2X cameras reconfig to nominal ops [Complete]
EVA Photo Downlink [Deferred]
Set up the first Orlan battery pack 825М3 for discharge and start discharging the first pack Report to MCC the number of the installed unit [Complete]
US Tools Stowage after EVA-37A [Complete]
Drying Orlan-MK Water Feed Line [Complete]
Psychological Evaluation Program (WinSCAT) [Complete]
Periodic Fitness Evaluation - Nom Ops [Complete]
AQM - Sieve Cartridges Checkout [Complete]
Orlan-MK Drying - start [Complete]
MLC Reboot [Complete]
ENERGY - Urine Sample Collection [Complete]
ENERGY - Water Sample Collection [Complete]
ТКГ 420/DC1 Interface Leak Check. [Complete]
ST-ISS - Crew gathers materials, sets up camcorder, reads the book [Complete]
ST-HIST - Documentation Photo [Complete]
Pille Dosimeter Reading after RS EVA [In Work]
EVA Video Data Downlink from Glisser-M [Complete]
СОЖ Maintenance [Complete]
TOCA Data Recording [Complete]
DC1-СУ Transfer Hatch Opening after EVA-37A [Complete]
СУ-ТКГ420 Hatch Opening after EVA-37A [Complete]
Orlan-MK Drying (switchover to a backup fan) [Complete]
ТКГ 420 (DC1) deactivation, air duct installation [Complete]
Exercise data downlink [Complete]
ENERGY Diet Log of Lunch [Complete]
Water Container (EDV) Remove and Replace [Complete]
ISERV - Experiment Procedure Review [Complete]
MATRYOSHKA-R. Monitoring Lulin-5 readings [Complete]
WHC Fill [Complete]
Food Warmer Troubleshooting Part 2 [Complete]
ISERV - Crew PL Conference [Complete]
Microbial Monitoring of the Environment (media 1 results obtained from locations Nos. 1-8) [Complete]
Orlan-MK Drying - terminate [Complete]
Setting up Orlan and БСС for storage mode [Complete]
ENERGY Diet Log of Dinner
Task List

ARED Cylinder Evac
IMS Update
National Geographic B-roll part 4 [In Work]
KTO Replace
ЕДВ-У R&R
EVA Video Data Downlink from Glisser-M
Preparation of reports for Roskosmos site
URAGAN. Observation and photography
EKON-M. Observations and photography
A/L PCS Troubleshoot [Complete]
EVA Wanted Poster Print [Complete]
HMS Trash Gather [Complete]
Spotlight Audit [Complete]
St Moon Read. [Complete]
Ground

Nominal Commanding
Columbus Cycle14 transition [Complete]
Three-Day Look Ahead:

Wednesday, 01/29: Capillary Flow Experiments, JEM LTL Pump Troubleshooting Part 2
Thursday, 01/30: CIR Manifold Bottle R&R, ISERV Pointing Mount repair, PEPS Audit
Friday, 01/31: Cubsat Deployer Install
QUICK ISS Status - Environmental Control Group:

Component - Status
Elektron - On
Vozdukh - Manual
[СКВ] 1 - SM Air Conditioner System ("SKV1") - On
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Off
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Process
Urine Processing Assembly (UPA) - Normal
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full up

Please follow SpaceRef on Twitter and Like us on Facebook.